OpenStudy (haleyelizabeth2017):

Which double-angle or half-angle identity would you use to verify that 1+cos 2α=(2)/(1+tan 2α)

OpenStudy (solomonzelman):

(I will use x instead, if you don't mind) Your question is like this? \(\large\color{#000000 }{ \displaystyle 1+\cos(2x) =\frac{2}{1+\tan(2x)} }\)

OpenStudy (haleyelizabeth2017):

Sorry

OpenStudy (haleyelizabeth2017):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @SolomonZelman (I will use x instead, if you don't mind) Your question is like this? \(\large\color{#000000 }{ \displaystyle 1+\cos(2x) =\frac{2}{1+\tan^2x} }\) \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (solomonzelman):

Are you allowed to play with both sides or you can only use one side?

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle 1+\cos(2x) =\frac{2}{1+\tan^2x} }\) \(\large\color{#000000 }{ \displaystyle 1+ \cos(2x) =\frac{2}{\sec^2x} }\) \(\large\color{#000000 }{ \displaystyle 1+ \cos(2x) = 2\cos^2x }\) \(\large\color{#000000 }{ \displaystyle 1+ \cos(2x) = 2\cos^2x -1+1 }\) \(\large\color{#000000 }{ \displaystyle 1+ \cos(2x) = \cos^2x+\cos^2x -1+1 }\) \(\large\color{#000000 }{ \displaystyle 1+ \cos(2x) = \cos^2x -\sin^2x+1 }\) this is how I would go about solving the problem

OpenStudy (solomonzelman):

you are using \(\cos(2\theta)=\cos^2\theta-\sin^2\theta=2\cos^2\theta-1 \) but, you are using it backwards

OpenStudy (solomonzelman):

you should be able to somple the problem without me...

OpenStudy (haleyelizabeth2017):

So sorry, I had a customer walk in.